Michael is updating his estate plan for himself and his family. He would like to provide an income of $ 3000 every month starting 101 2 years from now and continuing for the next 20 years. He has started his account with an initial deposit of $ 10,000 and he knows his life insurance, maturing in five years, will have a cash value of $ 150,000. To make up the difference , Michael has decided to make monthly deposits in the account. How much should each deposit be if all interest is computed at 6 percent compounded monthly ?
1
Expert's answer
2021-08-30T16:38:30-0400

It is assumed that, 10.5 years from now.


present value of future withdrawals as on 10.5 years

"=3,000\times(\frac{1-(1+0.5\%)^{-114})}{0.5\%})"


"=3,000\times 86.73416\\=\$260,202.5"


monthly payment

"=\frac{260202.5-(150,000\times(1+0.5\%)^{66}-(10,000\times(1+0.5\%)^{126}))}{(\frac{((1+0.5\%)^{126}-1)}{0.5\%}}"


"=\frac{32,982.12}{174.9331}"


"=\$188.5413" or

"=\$188.54"


Therefore, the monthly payment is $188.54